OpenStudy (anonymous):

A baby mouse weighs approximately 0.0022 pounds at birth. Write this weight in scientific notation.

OpenStudy (anonymous):

2.2* -10^3

OpenStudy (anonymous):

is that right.. because you move 3 decimals to the right.. so its negative...?

OpenStudy (anonymous):

i think so?

OpenStudy (anonymous):

Look at this...... :D \[2.2*\frac{1}{1000}=2.2*\frac{1}{10^{3}}=2.2*10^{-3}\]